# with Instruction Back-Translation

### 12 Jan 2024 v1

@@ Abstract
In this paper, we introduce KunThe dataset is named Kun as Chinese pronunciation of Humpback Li et al. (2023c)., a novel approach for creating high-quality instruction-tuning datasets for large language models (LLMs) without relying on manual annotations. Adapting a self-training algorithm based on instruction back-translation and answer polishment, Kun leverages unlabelled data from diverse sources such as Wudao, Wanjuan, and SkyPile to generate a substantial dataset of over a million Chinese instructional data points. This approach significantly deviates from traditional methods by using a self-curation process to refine and select the most effective instruction-output pairs. Our experiments with the 6B-parameter Yi model across various benchmarks demonstrate Kun’s robustness and scalability. Our method’s core contributions lie in its algorithmic advancement, which enhances data retention and clarity, and its innovative data generation approach that substantially reduces the reliance on costly and time-consuming manual annotations. This methodology presents a scalable and efficient solution for improving the instruction-following capabilities of LLMs, with significant implications for their application across diverse fields. The code and dataset can be found at https://github.com/Zheng0428/COIG-Kun
@@ 1 Introduction
The development of large language models (LLMs) relies on human-annotated datasets, a method facing scalability issues due to the significant resources required. Our study introduces Kun, a novel solution utilizing unlabelled data to create a high-quality instruction-tuning dataset. This method diverges from manual annotations, employing a self-training algorithm that includes a unique process called AP (Answer Polishment),
@@
We opt for the 6B-parameter Yi model due to its open-source nature and dependable performancehttps://github.com/01-ai/Yi. Its efficacy is tested and proven across diverse dataset sizes, including widely recognized benchmarks like C-EVAL and CMMLU. A comprehensive human evaluation uses 500 prompts from ShareGPT-zh, covering various tasks. Responses generated by our model are compared with those from other models, showcasing the superiority of our Kun-52k variant. Additionally, we evaluate the quality of our dataset, which includes 1,000 instruction-output pairs each from sources like Wudao (Xue et al., 2022), Wanjuan (He et al., 2023), and SkyPile (Wei et al., 2023). This evaluation, focusing on clarity, feasibility, practicality, and alignment, ensures the high quality of our dataset.
The key contributions of our work are:
@@
  • Data Creation: Over a million diverse Chinese instructional data points are produced from sources like Wudao, Wanjuan, and SkyPile, surpassing traditional crowdsourced annotations in quality and reducing reliance on manual annotation.
@@ 2 Related Work
Instruction Tuning
Instruction tuning is widely recognized as a key technique for activating LLMs to adhere to human conversational norms.  (Mishra et al., 2022; Wang et al., 2022b, 2023b). Instruction tuning empowers various domain-specific or task-specific LLMs, including natural language generation evaluation (Jiang et al., 2023), math (Yue et al., 2023; Xu et al., 2023a; Azerbayev et al., 2023), code Luo et al. (2023), music Li et al. (2023a); Deng et al. (2023), and medicine Wang et al. (2023a). Instruction tuning not only tailors the models’ task-specific responsiveness but also bolsters their cross-task generalization capabilities, thus enhancing performance across various dynamic application contexts (Wei et al., 2021; Sanh et al., 2022; Mishra et al., 2022; Wang et al., 2022b). Recent studies have broadened the scope of instruction tuning to encompass a wider array of general tasks, notably incorporating input from users of language models (Ouyang et al., 2022; Peng et al., 2023).
However, the open-source community is still lacking high-quality Chinese instruction tuning corpora. Current datasets, like Firefly (Yang, 2023), COIG (Zhang et al., 2023), BELLE (Ji et al., 2023), MOSS (Sun et al., 2023), and OL-CC (OL-CC, 2023), face issues such as limited scope, poor quality, commercial restrictions, or insufficient coverage. This gap hampers the advancement of LLMs in effectively processing and executing complex Chinese instructions, highlighting the critical need for more diverse and superior-quality datasets.
@@
Nevertheless, these instruction generation methodologies invariably require a robust teacher, whether a human annotator or a superior LLM. The ultimate potential of the model is limited by the teacher’s expertise or resource expenditure (Li et al., 2023c). To overcome this limitation, The SPIN (Chen et al., 2024) framework incorporates a self-play mechanism, enabling the Language Model to enhance its capabilities by engaging in battles against its instances. It generates training data from previous iterations, refining its strategy by distinguishing between responses generated autonomously and those derived from human-annotated data. This gradual process elevates the LLM from a nascent model to a robust one. Considering the vast amount of knowledge present in web text, Humpback introduces a technique based on Instruction Backtranslation. This method allows a base model to independently utilize vast amounts of unlabeled data to generate a high-quality instruction tuning dataset. However, empirical findings indicate that the effectiveness of this method is still constrained by the seed model’s performance and its ability to discern truly high-quality data.
@@ 3 Method
Our training approach requires a basic model, high-quality seed data, and a large amount of primary data. This primary data mainly includes a large volume of unlabeled data, unlike labeled data, which comes with predefined instructions and is ready for training. The main source of this data is web corpora, which contain various documents on human-interest topics, written by individuals and are crucial for several reasons:
  • Content Diversity: They encompass various subjects, mirroring the extensive scope of human interests.
  • Absence of Predefined Outputs: Unlike labeled data, these documents lack predetermined outcomes or objectives.
  • Potential for High-Quality Data: Certain subsets of these documents may be particularly suitable for generating text based on specific user instructions.
With these attributes, we can efficiently extract and use high-quality data from these rich corpora, adding specific instructions. Our innovative method improves language model training through a new data augmentation technique. Based on self-alignment principles, it involves carefully choosing and refining instructional data to fine-tune language models, as shown in Figure 1. This method allows us to gather a large amount of quality labeled data at a low cost, avoiding the expensive and time-consuming process of manual labeling.
Our method consists of two main steps:
@@ 3.1 Initialization
Seed data. We use 26,000 carefully chosen pairs of instructions and their corresponding outputs as our seed data. Each pair is hand-annotated and undergoes strict quality control to ensure high accuracy and consistency.
@@ 3.2 Generating Candidate Instructions with the Label Model
For each example ${y}_{i}$ in the unlabeled set, we utilize the backward model to infer a candidate instruction, denoted as $\hat{x}_{i}$. This procedure generates a collection of potential enhanced pairing data, represented as ($\hat{x}_{i}$, ${y}_{i}$). During annotation, initial filtering is based on perplexity (ppl) and length, excluding any data exceeding 512 tokens. We also discard data unsuitable for instructional use, like purely descriptive statements, while retaining useful data, such as commands and questions. We apply a specific prompt in this selection, keeping only data that triggers a positive response. The final labeled dataset contains instruction and output components, with instructions from the label model and outputs from primary data, known as candidate labeled data. Figure 5 shows the filter prompt used in this stage.
@@ 3.3 Refinement of Candidate Labeled Data Using the Primary Chat Model
  • Focused Scoring of Instruction Component: This technique only assesses the instruction part (output from the label model). High-scoring instructions are chosen, and then the output part of these selected data is refined."
Our analysis shows that the second method is more effective than the first. In the first method, good outputs are often discarded because of poor instructions from the label model, and the reverse is also true, causing unnecessary exclusions. Moreover, this approach sometimes retains data with one poor quality component because the other component is high quality, leading to uneven data quality and negatively impacting further training.
In contrast, the second method only scores the instruction component, effectively removing irrelevant data and favoring quality instructions, which improves data retention. Yet, it doesn’t assess the output, sometimes leading to suitable instructions paired with unsuitable outputs. To address this, we use the primary chat model to evaluate and refine the instructions and outputs, ensuring they align well. This approach produces high-quality labeled data. The score and refine prompts we used in this process are shown in Figure 6.
@@ 4.1.2 Baselines
For the Kun dataset, we annotated command data from three sources: Wudao, Wanjuan, and Skypile. This process has enriched the dataset, now made open-source and available for community usehttps://huggingface.co/datasets/m-a-p/COIG-Kun. Quantitative details of this augmented dataset are provided in Figure 7.In evaluating the performance of Kun, our study contrasts it with data curated from four prominent Chinese open-sourced datasets. These datasets are Firefly (Yang, 2023), COIG (Zhang et al., 2023; BAAI, 2023a, b), OL-CC (OL-CC, 2023), and BELLE (Ji et al., 2023), each unique in its composition and focus, providing a comprehensive basis for comparison.
@@ 4.1.3 Evaluation
Human Evaluation. To assess the general quality of model responses, we conduct human evaluations using a test set of 500 prompts sampled from ShareGPT-zh. These prompts, derived from actual user inquiries, encompass a diverse array of tasks, such as creative writing, information seeking, providing guidance, logical reasoning, storytelling, problem-solving, etc.
For the evaluation, responses generated by different models for each prompt are presented side-by-side. Human evaluators are asked to choose their preferred answer, providing a direct comparison of model performance. In total, eight models were compared, including three variants of the Kun model (fine-tuned with different sizes of our curated dataset), four baseline models (fine-tuned with public datasets), and one mixed model (fine-tuned with a combination of Kun and baseline datasets). This resulted in 28 pairwise comparisons across all test prompts. For this evaluation, we engage a team of experienced crowdsource annotators, ensuring a balanced and unbiased assessment. Detailed examples that show the comparison process can be found in Figure 12.
Standard Benchmarks. In addition to human evaluations, the models are also assessed using two standard benchmarks for Chinese LLM evaluation: C-EVAL  (Huang et al., 2023) and CMMLU  (Li et al., 2023b).
@@
Figure 7: Distribution of Data Sources
Figure 8: Human preference evaluation, comparing Kun-52k with other models.
Figure 8: Human preference evaluation, comparing Kun-52k with other models.
Figure 9: Human preference evaluation, comparing the mix model with other models.
@@ 4.2.1 Statistics and Diversity
Our work involve the purification of approximately 370,000 high-quality instruction-output pairs from the Wudao, Wanjuan, and Skypile datasets. This substantial dataset provide a diverse and rich foundation for our study.
@@ 4.2.2 Quality Evaluation
To ensure objectivity and reliability, three different evaluators evaluate each instruction-output pair. We conducted a detailed comparison of the options assessed by different evaluators, and the results demonstrate a significant degree of consistency in their judgments. Further details on evaluating identical ratings are presented in Figure 10 and Table 2. Examples that demonstrate the process of assessing can be found in Figure 12
4.3 Result Analysis
@@ 4.3.1 Human Evaluation
Table 3: Performance Statistics on Standard Benchmarks.The best results in each section are Bold , the second-best results are underlined, while the results of our best model are in Blue.
Our human evaluation results are illustrated through the heatmap in Figure 11, which shows the model vs. model win rates, with color variations indicating the relative performance strengths. The heatmap highlights that the Kun-360k model emerges as the most dominant, followed by the mixed model, showcasing their superior ability to handle a wide range of prompts. In contrast, the baseline models Firefly-52k, COIG-38k, and Belle-52k, garner lower preference percentages. This suggests that despite their strengths, these models may not align as closely with user expectations or prompt requirements as the Kun models.
Further analysis is provided in Figures 9 and 9, which depict the human preference evaluation win rates of Kun-52k and the mixed model, respectively, when compared with other models. Notably, the mixed model demonstrates a higher win rate than Kun-52k, indicating its enhanced effectiveness in meeting human evaluators’ preferences.
Figure 11: Heatmap of Model Comparative Win Rates in Human Evaluations
4.3.2 Standard Benchmarks
The performance of the models on standard benchmarks, specifically C-EVAL and CMMLU, provides an objective measure of their capabilities in understanding and reasoning within the Chinese context. These benchmarks assess the model’s ppl and generation quality (gen), key indicators of their language comprehension and output proficiency. Table 3 presents the performance statistics of various models on the C-EVAL and CMMLU benchmarks. As shown, we evaluate numerous models, including different sizes of the Kun model, baseline models, and mixed models. Each model’s performance is measured in terms of perplexity and generation quality, providing a comprehensive view of its strengths and weaknesses.
From the table, we observe that Kun-360k generally exhibits lower perplexity and higher generation quality, confirming its top-tier performance in language understanding and generation. Interestingly, the mixed model and Firefly display robust performance, with the mixed model often outperforming Kun-52k. The baseline models and smaller Kun variants present mixed results, excelling in some metrics while falling short in others. These highlight potential areas for further improvement in model training and fine-tuning strategies.
@@ 5 Conclusion
Our approach represents a breakthrough in instruction-tuning for LLMs, utilizing a novel self-training algorithm to leverage over a million quality Chinese instructional data points from diverse sources effectively. This strategy, different from manual annotations, not only enhances the instruction-following capabilities of LLMs but also ensures the high quality and diversity of training data. Empirical evaluations using the 6B-parameter Yi model across benchmarks like C-EVAL, CMMLU, and human evaluations, have demonstrated its robustness and scalability, particularly of the Kun-52k variant. Innovations within our approach, such as AP, have notably improved data retention and clarity, offering a scalable, efficient way to augment LLMs’ instructional capabilities. This research not only progresses the field but also broadens LLMs’ application scope, offering a novel solution to the challenges in developing instruction-tuning datasets.
Appendix A Example Appendix
This is a section in the appendix.

### 23 Feb 2024 v2

@@ Abstract
In this paper, we introduce KunThe dataset is named Kun as Chinese pronunciation of Humpback Li et al. (2023c)., a novel approach for creating high-quality instruction-tuning datasets for large language models (LLMs) without relying on manual annotations. Adapting a self-training algorithm based on instruction back-translation and answer polishment, Kun leverages unlabelled data from diverse sources such as Wudao, Wanjuan, and SkyPile to generate a substantial dataset of over a million Chinese instructional data points. This approach presents a novel departure from traditional methods by using a self-curation process to refine and select the most effective instruction-output pairs. Our experiments with the 6B-parameter Yi model across various benchmarks demonstrate Kun’s robustness and scalability. Our method’s core contributions lie in its algorithmic advancement, which enhances data retention and clarity, and its innovative data generation approach that substantially reduces the reliance on costly and time-consuming manual annotations. This methodology presents a scalable and efficient solution for improving the instruction-following capabilities of LLMs, with significant implications for their application across diverse fields. The code and dataset can be found at https://github.com/Zheng0428/COIG-Kun
Kun: Answer Polishment for Chinese Self-Alignment
with Instruction Back-Translation
[…]
@@ 1 Introduction
The development of large language models (LLMs) relies on human-annotated datasets, yet the creation of such datasets typically faces scalability issues due to the significant resources required. Our study introduces Kun, a novel approach leveraging unlabelled data to create a high-quality instruction-tuning dataset. This method diverges from manual annotations, employing a self-training algorithm that includes a unique process called AP (Answer Polishment),
@@
We opt for the 6B-parameter Yi model due to its open-source nature and dependable performancehttps://github.com/01-ai/Yi. Its efficacy is tested and proven across diverse dataset sizes, including widely recognized benchmarks like C-EVAL (Huang et al., 2023) and CMMLU (Li et al., 2023b). To evaluate the performance of the model, we design acomprehensive human evaluation which contains 500 prompts from ShareGPT-zh, covering various tasks. Responses generated by our model are compared with those from other models, showcasing the superiority of our Kun-52k variant. Further details can be found in 4.3.1. Additionally, we evaluate the quality of our dataset, which includes 1,000 instruction-output pairs each from sources like Wudao (Xue et al., 2022), Wanjuan (He et al., 2023), and SkyPile (Wei et al., 2023). This evaluation, focusing on clarity, feasibility, practicality, and alignment, ensures the high quality of our dataset. The key contributions of our work are:
@@
  • Large-scale high quality data creation: Over a million diverse Chinese instructional data points are produced from sources like Wudao, Wanjuan, and SkyPile, surpassing traditional crowdsourced annotations in quality and reducing reliance on manual annotation.
@@ 2 Related Work
Instruction Tuning Instruction tuning is widely recognized as a key technique for activating LLMs to adhere to human conversational norms.  (Mishra et al., 2022; Wang et al., 2022b, 2023b). Instruction tuning empowers various domain-specific or task-specific LLMs, including natural language generation evaluation (Jiang et al., 2023), math (Yue et al., 2023; Xu et al., 2023a; Azerbayev et al., 2023), code Luo et al. (2023), music Li et al. (2023a); Deng et al. (2023), and medicine Wang et al. (2023a). Instruction tuning not only tailors the models’ task-specific responsiveness but also bolsters their cross-task generalization capabilities, thus enhancing performance across various dynamic application contexts (Wei et al., 2021; Sanh et al., 2022; Mishra et al., 2022; Wang et al., 2022b). Recent studies have broadened the scope of instruction tuning to encompass a wider array of general tasks, notably incorporating input from users of language models (Ouyang et al., 2022; Peng et al., 2023).
However, the open-source community is still lacking high-quality Chinese instruction tuning corpora. Current datasets, like COIG (Zhang et al., 2023), BELLE (Ji et al., 2023), MOSS (Sun et al., 2023), and OL-CC (OL-CC, 2023), face issues such as limited scope, poor quality, commercial restrictions, or insufficient coverage. This gap hampers the advancement of LLMs in effectively processing and executing complex Chinese instructions, highlighting the critical need for more diverse and superior-quality datasets.
@@
Nevertheless, these instruction generation methodologies all require a robust teacher.The ultimate potential of the model is limited by the teacher’s expertise or resource expenditure (Li et al., 2023c). To overcome this limitation, the SPIN (Chen et al., 2024) framework incorporates a self-play mechanism, It generates training data from previous iterations, refining its strategy by distinguishing between responses generated autonomously and those derived from human-annotated data. This gradual process elevates the LLM from a nascent model to a robust one. Considering the vast amount of knowledge present in web text, Humpback Li et al. (2023c) introduces a technique based on Instruction Backtranslation. This method allows a base model to independently utilize vast amounts of unlabeled data to generate a high-quality instruction tuning dataset. However, empirical findings indicate that the effectiveness of this method is still constrained by the seed model’s performance and its ability to discern truly high-quality data.
@@ 3 Method
Our training methodology necessitates a foundational model, high-quality seed instruction data, and a substantial volume of unlabeled data, with the primary source being web text. Given the extensive content diversity inherent in large-scale web documents, which encompass a wide array of topics such as music, art, technology, etc., reflecting the broad spectrum of human interests, certain subsets of these documents may be more apt for generating instructions. Unlike labeled data, these documents lack predetermined outcomes or objectives. This method involves the refinement and optimization of data selection during the fine-tuning process of Large Language Models (LLMs), as illustrated in Figure 1. This approach allows for the collection of a significant volume of instructional data at a low cost, circumventing the exorbitant expenses associated with manual labor, in a manner that is both academically rigorous and professional. Our method consists of two main steps:
@@ 3.1 Initialization
Seed data. We use 26,000 instructions and their corresponding outputs as our seed data. Each pair is hand-annotated and undergoes strict quality control to ensure high accuracy and consistency.
@@ 3.2 Generating Candidate Instructions with the Label Model
For each example ${y}_{i}$ in the unlabeled set, we utilize the backward model to infer a candidate instruction, denoted as $\hat{x}_{i}$. This procedure generates a collection of potential enhanced pairing data, represented as ($\hat{x}_{i}$, ${y}_{i}$). During annotation, initial filtering is based on perplexity (ppl) and length, and excludes any data exceeding 512 tokens. We also discard data unsuitable for instructional use, like purely descriptive statements, while retaining useful data, such as commands and questions. We apply a filter prompt in this selection, keeping only data that triggers a positive response. The final labeled dataset contains instruction and output components, with instructions from the label model and outputs from primary data, known as candidate labeled data. Figure 4 shows the filter prompt used in this stage.
@@ 3.3 Refinement of Candidate Labeled Data Using the Primary Chat Model
  • Focused Scoring of Instruction Component: This technique only assesses the instruction part (output from the label model). High-scoring instructions are chosen, and then the output part of these selected data is refined.
Our analysis shows that the second method is more effective than the first. In the first method, good outputs are often discarded because of poor instructions from the label model, and the reverse is also true, causing unnecessary exclusions. Moreover, this approach occasionally retains data with one poor quality instruction because the corresponding output is high quality, and vice versa, leading to uneven data quality and negatively impacting further training.
In contrast, the second method only scores the instruction component, as in instruction tuning for LLM, instructions are often considered more important than outputs,Yet, it doesn’t assess the output, sometimes leading to suitable instructions paired with unsuitable outputs. To address this, we use the primary chat model to evaluate and refine the instructions and outputs, ensuring they align well. This approach produces high-quality labeled data. The score and refine prompts we used in this process are shown in Figure 5.
@@ 4.1.2 Baselines
For the Kun datasethttps://huggingface.co/datasets/m-a-p/COIG-Kun, we annotated command data from three sources: Wudao, Wanjuan, and Skypile. Quantitative details of this augmented dataset are provided in Figure 6. In evaluating the performance of Kun, our study contrasts it with data curated from four prominent Chinese open-sourced datasets, including COIG (Zhang et al., 2023; BAAI, 2023a, b), OL-CC (OL-CC, 2023), and BELLE (Ji et al., 2023). These datasets are unique in their composition and focus, providing a comprehensive basis for comparison.
@@ 4.1.3 Evaluation
Human Evaluation. To assess the general quality of model responses, we conduct human evaluations using a test set of 500 prompts sampled from ShareGPT-zh. These prompts, derived from real world user inquiries, encompass a diverse array of tasks, such as creative writing, information seeking, providing guidance, logical reasoning, storytelling, problem-solving, etc.
For the evaluation, responses generated by different models for each prompt are presented side-by-side. Human evaluators are asked to choose their preferred answer, providing a direct comparison of model performance. In total, eight models were compared.For this evaluation, we engage a team of experienced crowdsource annotators, ensuring a balanced and unbiased assessment. Detailed examples that show the comparison process can be found in Figure 10. Standard Benchmarks. In addition to human evaluations, the models are also assessed using two standard benchmarks for Chinese LLM evaluation: C-EVAL  (Huang et al., 2023) and CMMLU  (Li et al., 2023b).
@@ 4.2.1 Statistics and Diversity
Our work involve the purification of approximately 377,592 high-quality instruction-output pairs from the Wudao, Wanjuan, and Skypile datasets. This substantial dataset provide a diverse and rich foundation for our study.
@@ 4.2.2 Quality Evaluation
To ensure objectivity and reliability, three different evaluators evaluate each instruction-output pair. The consistency rates for the evaluation across the three dimensions of the instructions have all exceeded 90%, and the evaluation of the instruction-response also demonstrates a high rate of consistency. This results demonstrate a significant degree of consistency in their judgments. Further details on evaluating identical ratings are presented in Figure 7 and Table 2. Examples that demonstrate the process of assessing can be found in Figure 10
4.3 Result Analysis
@@ 4.3.1 Human Evaluation
Our human evaluation results are illustrated through the heatmap in Figure 8, which shows the model vs. model win rates, with color variations indicating the relative performance strengths. The heatmap highlights that the Kun-52k model emerges as the most dominant, followed by the mixed model, showcasing their superior ability to handle a wide range of prompts. In contrast, the baseline models COIG-39k, and Belle-52k, garner lower preference percentages. This suggests that despite their strengths, these models may not align as closely with user expectations or prompt requirements as the Kun models. Further analysis is provided in Appendix A.4.
Figure 8: Heatmap of Model Comparative Win Rates in Human Evaluations
4.3.2 Standard Benchmarks
The performance of the models on standard benchmarks, Table 3 presents the performance statistics of various models on the C-EVAL and CMMLU benchmarks. As shown, we evaluate numerous models, including different sizes of the Kun model, baseline models, and mixed models. Each model’s performance is measured in terms of perplexity and generation quality, providing a comprehensive view of its strengths and weaknesses.
From the table, we observe that Kun-39k generally exhibits lower perplexity and higher generation quality, confirming its top-tier performance in language understanding and generation. Interestingly, the mixed model display robust performance, with the mixed model often outperforming Kun-52k. The baseline models and smaller Kun variants present mixed results, excelling in some metrics while falling short in others. These highlight potential areas for further improvement in model training and fine-tuning strategies.
@@ 4.3.2 Standard Benchmarks
Table 3: Performance Statistics on Standard Benchmarks.The best results in each section are Bold , the second-best results are underlined, while the results of our best model are in Blue. PPL:Perplexity GEN:Generation
@@ 5 Conclusion
Our approach represents a breakthrough in instruction-tuning for LLMs, utilizing a novel self-training algorithm to leverage over a million quality Chinese instructional data points from diverse sources effectively. This strategy, different from manual annotations, not only enhances the instruction-following capabilities of LLMs but also ensures the high quality and diversity of training data. Empirical evaluations using the 6B-parameter Yi model across benchmarks like C-EVAL, CMMLU, and human evaluations, have demonstrated its robustness and scalability. Innovations within our approach, such as AP, have notably improved data retention and clarity, offering a scalable, efficient way to augment LLMs’ instructional capabilities. This research not only progresses the field but also broadens LLMs’ application scope, offering a novel solution to the challenges in developing instruction-tuning datasets.
6 Ethics Statement
This study adheres to the highest ethical standards, ensuring all research activities are conducted with a commitment to responsibility and respect for participant rights. Our ethical policy encompasses data usage, intellectual property rights respect, and research transparency. To safeguard data privacy and security, particularly when handling unlabeled data from sources like Wudao, Wanjuan, and SkyPile, we implement stringent measures to comply with data protection laws, especially concerning personal information. This involves anonymizing and desensitizing data prior to utilization. In terms of intellectual property rights, we ensure that all employed data and generated guiding points adhere to applicable copyright laws and intellectual property agreements, thereby avoiding infringement on any third-party intellectual property rights. Moreover, we pledge to provide a comprehensive account of our research methodology, detailing the processes of data generation, model training, and evaluation, to facilitate reproducibility and validation by the academic community.
7 Limitations
This study introduces an innovative methodology for generating data to reduce the reliance on costly and time-consuming manual annotations. However, there are limitations impacting the generalizability and scalability of our findings. Firstly, the diversity of data: the methodology, while capable of generating instructional data from a vast pool of unlabeled data, may produce data whose quality and diversity are constrained by the original data source’s breadth and caliber. This is particularly relevant when generating instructions for niche or specialized domains, where additional methods might be necessary to guarantee comprehensive coverage and precision. Secondly, the generalization capability of the model: although tests on the 6B-parameter Yi model confirm the methodology’s efficacy, its performance and applicability could differ among models of various sizes and tasks. Its effectiveness might require further investigation, especially in smaller models or those not designed for Chinese language processing. Thirdly, the assumptions underlying the algorithm: the study’s self-training algorithm relies on instruction back-translation and answer embellishment. These premises might not hold across all instruction or answer types, notably for tasks demanding specialized knowledge or creativity, where the generated instructional data may not adequately reflect complex cognitive processes. Future Directions: Subsequent research should examine the methodology’s applicability to large-scale language models across different languages and the generation of high-quality instructional data in specialized fields such as medicine or law. Additionally, advancing self-training algorithms to more effectively manage intricate and specialized instructions represents a crucial avenue for exploration.
Appendix A Appendix